Circular No. 6257

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1993J IN NGC 3031 P. Garnavich, P. Challis, and R. Kirshner, Center for Astrophysics;

and L. Wells, University of Arizona, communicate: “The brightness of SN 1993J has remained nearly constant in BVRI over the past four months, after having faded during the first half of 1995 at rates of 0.002 mag/day in R and ≈ 0.003 mag/day in B. On Nov. 2, R = 18.20 and B = 19.27 (± 0.05). Spectra obtained by P. Berlind with the 1.5-m Tillinghast telescope at Mt. Hopkins on Oct. 27 show the continued presence of the strong, broad (halfwidth 10 000 km/s) Hα emission feature. The spectrum has not changed substantially since 1995 Apr., except that the profile of the Hα emission has become more symmetric. A narrow (unresolved) Hα emission line is superimposed on the broad feature at a wavelength of 656.03 nm.”

COMET 122P/DE VICO C. C. Petersen and J. C. Brandt write: “Results from the ULYSSES

mission have clearly established that the equatorial and polar comets are exposed to a very different solar-wind environment. Comet 122P/de Vico is an excellent target for purposes of recording the variation of cometary properties with heliographic latitude and direct comparison with in situ measurements obtained by the ULYSSES spacecraft. 122P has shown good plasma activity, including a plasma tail and at least one disconnection event; it has a very-high-inclination orbit, which covers a large latitude range in 1995 Oct. and Nov., and it has approximately the same latitude as ULYSSES during Nov. On behalf of the Ulysses Comet Watch, we request coverage of 122P/de Vico with images that show plasma structures; please communicate such images to us at LASP, CB 392, University of Colorado, Boulder, CO 80309, U.S.A. (telefax 303-492-6946).”

COMET C/1995 Q1 (BRADFIELD) Total visual magnitude estimates (IAUC 6251): Oct. 25.09 UT, 10.0 (L.

Kiss, Szeged, Hungary, 0.20-m reflector); 26.50, 9.5 (A. Hale, Cloudcroft, NM, 10×50 binoculars); 28.49, 8.3 (R. Keen, Mt. Thorodin, CO, 0.15-m reflector); 29.50, 9.2 (C. E. Spratt, Victoria, BC, 0.20-m reflector); Nov. 4.18, 7.5 (A. Diepvens, Balen, Belgium, 0.15-m refractor); 5.21, 8.9 (J. Carvajal, Madrid, Spain, 0.21-m reflector).

1995 November 7 Daniel W. E. Green